OpenStudy (anonymous):

PLEASE HELP ME Solve this system of equations using any method. 3x - y = 4 x + 2y = 13 A. (-1, 7) B. (4, 8) C. (5, 11) D. (3, 5)

OpenStudy (mathmath333):

by any method make any one coefficient same of both the equations example by multiplying eqn 1 by 2

OpenStudy (anonymous):

i got B

OpenStudy (mathmath333):

no thats wrong

OpenStudy (anonymous):

D

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ 3x-y=4\hspace{.33em}\\~\\~\\ \implies 6x-2y=8\hspace{.33em}\\~\\ x + 2y = 13\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

add the last two equations

OpenStudy (anonymous):

D?

OpenStudy (mathmath333):

yes thats correct

OpenStudy (anonymous):

Now you are right

OpenStudy (anonymous):

thanks